When a certain amount of ethylene was burnt 6226kj heat was evolved. If heat of combustion of ethylene is 1411 kj then the volume of Oxygen at NTP that entered in reaction is ??

Grade:9

1 Answers

mark
6 years ago
c2h4+3o2------>2co2+2h2o
moles of o2 reacted=(6226/1411)*3
vol. of o2 reacted at STP=(6226/1411)*3*22.4=296.5 ltrs